# GREEN-CODE: Optimizing Energy Efficiency in Large Language Models for Code Generation

### v1

@@ V Implementation and Deployment
As specific algorithm, we chose the Proximal Policy Optimization (PPO) [32] algorithm, an RL method that has demonstrated effectiveness across various tasks in different domains. PPO offers stability in training by using clipping approaches to avoid excessively large updates, leading to more stable training. We employed single environment and shallow network configurations. We trained the network until convergence was observed in both episodic and stepwise rewards. On average, it took 200k-500k steps before observing convergence. A detailed description of the hyperparameter setup and results are provided in the next section.
@@
For deployment, we implemented an endpoint that deploys the LLM alongside an RL agent. The endpoint adheres to the Huggingface Inference API, enabling our LLMs to be used in the Huggingface VSCode extension [^1] https://github.com/huggingface/llm-vscode. Consequently, our setup can be utilized similarly to GitHub Copilot or other tools. Furthermore, we also integrated GREEN-CODE with an existing VSCode extension from Huggingface. Using our deployed LLM endpoint and extended configurations of the VSCode extension, we are able to specify runtime LLM behavior (e.g., num_predict) as well as RL agent thresholds to decide trade-off between resource consumption and accuracy. We implemented our framework using Python 3.12 programming language. GREEN-CODE prototype implementation, deployment configuration scripts, and details, including libraries and dependencies used, are available in an open-source GitHub repository at https://anonymous.4open.science/r/GREEN-CODE-03CC.
@@ VI-E2 Performance Evaluation on PY150 Dataset
VI-F Overhead Analysis

### v2

# GREEN-CODE: Learning to Optimize Energy Efficiency in LLM-based Code Generation

@@ V Implementation and Deployment
As a specific algorithm, we chose the Proximal Policy Optimization (PPO) [32] algorithm, an RL method that has demonstrated effectiveness across various tasks in different domains. PPO offers stability in training by using clipping approaches to avoid excessively large updates, leading to more stable training. We employed single environment and shallow network configurations. We trained the network until convergence was observed in both episodic and stepwise rewards. On average, it took 200k-500k steps before observing convergence. A detailed description of the hyperparameter setup and results are provided in the next section.
@@
For deployment, we implemented an endpoint that deploys the LLM alongside an RL agent. Consequently, our setup can be utilized similarly to GitHub Copilot or other tools. Furthermore, we also integrated GREEN-CODE with an existing VSCode extension from Huggingface. The endpoint adheres to the Huggingface Inference API, enabling our LLMs to be used in the Huggingface VSCode extension [^1] https://github.com/huggingface/llm-vscode. Using our deployed LLM endpoint and extended configurations of the VSCode extension, we are able to specify runtime LLM behavior (e.g., num_predict) as well as RL agent thresholds to decide trade-off between resource consumption and accuracy. We implemented our framework using Python 3.12 programming language. GREEN-CODE prototype implementation, deployment configuration scripts, and details, including libraries and dependencies used, are available in an open-source GitHub repository at https://github.com/Large-scale-Sustainable-Computing-LSC/green-code.
@@ VI-E2 Performance Evaluation on PY150 Dataset
VI-F Sensitivity Analysis of Context Length
In the previous subsection, we presented the context settings of the results 0.2 (that is, 20% of the code file as input). Figure 12 illustrates the performance of Llama3.2 on JavaCorpus in different context lengths of 0.2, 0.3, 0.5, and 0.6, while varying the agent thresholds as described earlier. The results are shown for CodeBLEU and energy, as these metrics exhibit overall behavior similar to other indicators.
The findings demonstrate that GREEN-CODE overall performs similarly in all context settings. However, in some higher context settings, the difference to the full model is larger than in lower context settings. For example, in the context setting 0.2, CodeBLEU is 17% lower at $T=0.6$, but only 5.28% lower at $T=0.92$. In contrast, in the context of 0.5, these values increase to 27.2% and 8.9%, respectively, almost double the loss observed in the context of the lower context. Similar patterns are evident in the other context settings. Although there is a small variation in performance across different input sizes, its behavior in terms of energy savings remains similar. Consequently, we reported primary results for context setting 0.2.
(a) CodeBLEU
(b) Total Energy.
Figure 12: Results of Llama on JavaCorpus, with different GC agent thresholds $T$, input contexts and baselines.
VI-G Analyzing KV Cache Impact
The experiments described above were conducted without KV caching due to the native incompatibility of KV caching with early exits. This incompatibility arises because caches may be unavailable when a deeper exit follows a shallower one. Although this aspect was not our primary focus, we implemented a basic KV cache propagation method, as outlined in previous studies [17], to evaluate the impact of KV caching. Figure 13 presents the results, including various accuracy-related metrics and the average number of layers utilized per experiment. The findings indicate that our approach achieves accuracy comparable to other reported methods. Optimizing KV caching during early exits, such as through parallel cache computation, remains a promising avenue for future research.
(a) Score metrics
(b) Avg. #Layers.
Figure 13: Results of Llama on JavaCorpus with KV caching.
VI-H Overhead Analysis
@@ VIII Conclusions
Acknowlegments
The research presented in this paper has been partially funded through the following projects: High-Performance Integrated Quantum Computing (HPQC), Austrian Research Promotion Agency (FFG) Nr. 897481; Transprecise Edge Computing (Triton), Austrian Science Fund (FWF), DOI: 10.55776/P36870; and Trustworthy and Sustainable Code Offloading (Themis), FWF, DOI: 10.55776/PAT1668223.